Solve
log5 3^x = log25 9^(1-2x)

Answer by lwsshak3(11628)

log5 3^x = log25 9^(1-2x)
convert to log base 10
log(3^x)/log(5)=log(9^(1-2x))/log25
log25=log5^2=2log5
log(3^x)/log(5)=log(9^(1-2x))/2log5
2log(3^x)=log(9^(1-2x)
log(3^2x)=log(9^(1-2x)
(3^2x)=(9^(1-2x)
(3^2x)=(3^(2(1-2x))
3^2x=3^(2-4x)
2x=2-4x
6x=2
x=1/3